英語 (English)

発信型英語 (Communicative English) (総科 ((総理 b)2 年))

木里恭子 · 非常勤講師 / 全学共通教育センター 2 単位後期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】The English style of writing differs from the Japanese style of writing. To be an effective writer in English language, students need to learn the English style of writing in addition to words and grammar rules. The purpose of taking this course is to learn the English method of direct development which will make students’ own writing not only clear and convincing, but interesting and readable.

【授業の概要】Students will learn the paragraph-writing skills in English. The course emphasizes both content and form: to help students express ideas and convey information in writing (1) with logical reasoning and adequate support and (2) with clarity of purpose, organization, and language. The course guides students through the composing process of prewriting, writing, and postwriting.

【キーワード】 EFL， Composition， Paragraphs， Rhetoric， Academic writing 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】The goal of this course is to help students improve their written

communication skills. By the end of the course, students will have an overview of academic writing. They will learn, practice, and master the basic composition skills and strategies.

【授業の計画】 1. Paragraph Organization

2. TS/ SS/ CS

3. Characteristics of Good Writing

4. Unity/ Coherence/ Cohesion

5. Narration

6. Narration

7. Narration

8. MIDTERM EXAM

9. Description

10. Description

11. Description 12. Persuasion

13. Persuasion

14. Persuasion

15. FINAL EXAM

16. Recapitulation 【教科書】 Blanchard, K and Root, C. Ready to Write 2: Perfecting Paragraphs (4th

ed.) 【参考書等】 [参考資料] 【成績評価の方法】 Portfolio Assignments = 60% Midterm & Final = 40% 【再試験の有無】 No make-up exams will be given. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221209 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 木里 . (オフィスアワー: Tue. (Preferred by appointment only.))

English

Communicative English (総科 ((総理 b)2 年))

Kyoko Kizato · Part-time Lecturer / Center for General Education 2 units 後期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ The English style of writing differs from the Japanese style of writing. To be an effective writer in English language, students need to learn the English style of writing in addition to words and grammar rules. The purpose of taking this course is to learn the English method of direct development which will make students’ own writing not only clear and convincing, but interesting and readable.

Outline⟩ Students will learn the paragraph-writing skills in English. The course emphasizes both content and form: to help students express ideas and convey information in writing (1) with logical reasoning and adequate support and (2) with clarity of purpose, organization, and language. The course guides students through the composing process of prewriting, writing, and postwriting.

�������⟩ EFL, Composition, Paragraphs, Rhetoric, Academic writing Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ The goal of this course is to help students improve their written

communication skills. By the end of the course, students will have an overview of academic writing. They will learn, practice, and master the basic composition skills and strategies.

Schedule⟩

1. Paragraph Organization

2. TS/ SS/ CS

3. Characteristics of Good Writing

4. Unity/ Coherence/ Cohesion

5. Narration

6. Narration

7. Narration

8. MIDTERM EXAM

9. Description

10. Description

11. Description 12. Persuasion

13. Persuasion

14. Persuasion

15. FINAL EXAM

16. Recapitulation Textbook⟩ Blanchard, K and Root, C. Ready to Write 2: Perfecting Paragraphs

(4th ed.) Reference⟩ [参考資料] Evaluation Criteria⟩ Portfolio Assignments = 60% Midterm & Final = 40% Re-evaluation⟩ No make-up exams will be given. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221209 Contact (Office-Hour, Room, E-mail)⟩

⇒ Kizato . (Office Hour: Tue. (Preferred by appointment only.))